TERAHERTZ ROTATIONAL SPECTROSCOPY OF THE SO RADICAL

\underline{M. A. MARTIN-DRUMEL}, A. CUISSET, S. ELIET, G. MOURET, F. HINDLE, Laboratoire de Physico-Chimie de l’Atmosphère, EA 4493, Université du Littoral Côte d’Opale, 59140 Dunkerque, France; O. PIRALI, Institut des Sciences Moléculaires d’Orsay, CNRS, UMR 8214, Université Paris XI, bât. 210, 91405 Orsay Cedex, France; SOLEIL Synchrotron, AILES beamline, L’orme des Merisiers, Saint-Aubin, 91192 Gif-Sur-Yvette, France.

We have recorded pure rotational transitions of SO in the THz spectral range using synchrotron-based Fourier-Transform (FT) FIR and continous wave (CW) THz techniques. A FT-FIR spectrum of SO has been recorded at the AILES beamline of SOLEIL synchrotron in the spectral range 44–93 cm^{−1} using a resolution of 0.001 cm^{−1} allowing an accuracy on line position of 0.00007 cm^{−1} (∼ 2 MHz). A multipass absorption discharge cell aligned to an absorption path length of 24 m has been used^{a}. A continuous electrical discharge (1 A / 980 V) in a flowing mixture of H_{2}S, He, H_{2} and air (respectively at pressure of 0.01, 1.15, 0.14 and 0.06 mbar) was used to produce SO. On this spectrum, 102 transitions of SO have been identified with N = 31 to 65. Among the observed lines, 99 are detected for the first time (22 new transitions belong to the HIFI spectral windows). Due to our limited instrumental resolution, transitions involving N ranging from 31 to 43 show unresolved fine structure triplets. Recently, in order to observe all fine structure components in the HIFI spectral windows, we have recorded a high resolution CW-THz spectrum of SO^{b}. At the time of the writing, this spectrum was under analysis.

aM. A. Martin-Drumel et al., Rev. Sci. Instrum. 82, 113106 (2011)

bS. Eliet et al., J. Mol. Struct. 1006, 13 (2011)